OpenStudy (anonymous):

Find the reference angle in radians for angle of 5pi/3. Exact Answer

OpenStudy (anonymous):

(5pi)/3 liek this?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hmm. exact answer... meaning no deimals?

OpenStudy (anonymous):

no exact answer meaning in radians and not degrees such as radical 3/2

OpenStudy (anonymous):

i'm stumped on how to get (5/3)pi into an exact number from what it already is given in radians... time to bring in some bigger guns

OpenStudy (anonymous):

@jdoe0001

OpenStudy (jdoe0001):

hmm the most you can do to that is expand by using the radian value for \(\bf \pi\) so \(\bf \cfrac{5\pi}{3}\qquad \pi \approx 3.1416\qquad thus \quad \cfrac{5\pi}{3}\implies \cfrac{5\cdot 3.1416}{3}\quad radians\)

OpenStudy (jdoe0001):

hmmm ohh... wait... you want the reference .. angle...dohh , ok

OpenStudy (jdoe0001):



OpenStudy (anonymous):

oh gosh, I didn't even notice that part of the question o.O haha

OpenStudy (anonymous):

It has to be in an Exact Answer though 5pi/3=360-300=60 degrees= pi/3

OpenStudy (jdoe0001):

yeap, you're correct, it's \(\bf \cfrac{\pi}{3}\)

OpenStudy (jdoe0001):

the \(\large \pi\) unit is already in radians value, so that's proper

OpenStudy (anonymous):

thank you for the help

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

though

OpenStudy (anonymous):

could you help me with Use the reference angle to find sec(11pi/6) lol

OpenStudy (jdoe0001):

 \(\bf sec\left(\cfrac{11\pi}{6}\right)\implies sec\left(\cfrac{\pi}{6}\right)\implies \cfrac{1}{cos\left(\frac{\pi}{6}\right)}\)